OpenStudy (anonymous):

Find the equation of the tangent plane to the graph of z=xy^2 at the point (1,1,1)

OpenStudy (anonymous):

Start with finding the gradient of your function.

OpenStudy (anonymous):

We haven't done that yet

OpenStudy (anonymous):

Okie doke, not a problem. Have you done partial derivatives yet?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay, first thing's first. Let's move that z over to the other side and let \[f(x) = xy^2 - z\]

OpenStudy (anonymous):

The gradiant of f is just going to be: (\[(f _{x}, f _{y},f _{z})\]

OpenStudy (anonymous):

so... \[(y ^{2},2yx,-1)\]

OpenStudy (anonymous):

Good, now plug in your point and this will give you your vector for the tangent plane.

OpenStudy (anonymous):

Just use (1,1,1)

OpenStudy (anonymous):

Oh, I see, yes, but your 2nd coordinate should be 2. 2(1)(1)

OpenStudy (anonymous):

my bad... (1,2,-1)

OpenStudy (anonymous):

Okay, this is the vector for the tangent plane. You've have a point on that plane and a vector. Where do we go from here?

OpenStudy (anonymous):

Some sort of linearization?

OpenStudy (anonymous):

Something like that. We know the general equation for this plane will have: <1,2,-1> So: x + 2y - z = d Plug your point in (1,1,1) in to that to get your d.

OpenStudy (anonymous):

So d=2

OpenStudy (anonymous):

Cool, so there's your equation: x + 2y - z = 2

OpenStudy (anonymous):

In general, every sort of problem like this uses these steps.

OpenStudy (anonymous):

That was actually really simple! Thank you!

OpenStudy (anonymous):

No problem, good luck. The hardest part about this level of calculus is just visualizing what your looking for. The math is pretty much just plug and chug. :D Good luck!

OpenStudy (anonymous):

Thanks!